Telescope.org
This website uses cookies. To read more about our use of cookies, click here.
To not be bothered by this message again, click the OK button.
OK

Current Status

We have recently carried out a servicing mission where the following tasks were completed: changed the PIRATE broad band filter set from Baader LRGB to Johnson BVR, changed the COAST camera to a 3k chip version of the camera already on PIRATE (FLI Proline), changed the COAST broad band filter set from Johnson BVR to Baader LRGB. The COAST focussing issue was resolved. The COAST Schmidt plate received a clean to remove built-up dust and debris as did the PIRATE primary mirror.

A further service mission has re-instated the Johnson filters on COAST and added UV and IR for PIRATE. There are some issues with the image size and calibration filters for COAST and we are also aware that PIRATE is currently out of focus and are working to resolve this.

COAST is currently unable to be used due to a router failure at the dome. PIRATE is working fine.

Planned Outages

None


Last updated: Monday 3 September 2018, 11:14



Back to the index page

© 2018 The Open University